1. Prove that ( 1 - cos2 x ) cosec2 x = 1 .

2. Prove that (tan x sin x ) + cos x = sec x .

3. Prove that Tan^4 x + Tan^2 x = sec^4 x - sec^2 x .

4. prove that Cot x + Tan x = sec x cosec x.

5. Prove that (1 + sin x) (1 - sin x) = (sec x - Tan x)2 .
1
Expert's answer
2020-08-11T19:10:10-0400

solution for question 2

(sinx/cosx .sinx) +cosx=

sinx/cosx +cosx=

(sin2x +cos2x)/cosx=

1/cosx

= secx

solution for question 1

(1 -cos2x)cosec2x=

(sin2x)cosec2x=

(sin2x) . 1/sin2x=

= 1


solution for question 4

cotx +tanx =

cosx/sinx + sinx/cosx=

(cos2x +sin2x)/(cosxsinx)=

1/(cosxsinx)=

1/cosx .1/sinx=

=secxcosecx


question 5 solution

(1 +sinx)(1 -sinx)

1 -sin2x =cos2x from the difference of two squares

(1/cosx +sinx/cosx)(1/cosx -sinx/cosx)=

(secx+Tanx)(secx -Tanx)

looking at the reverse method

(secx -Tanx)2 =(1 +sinx)(1 -sinx)

(1/cosx -sinx/cosx)

(1 -sinx)2/cos2x


question 3 solution

Tan4x + Tan2x = sec4x -sec2x

Tan2x(Tan2x +1)

recall cos2x + sin2x =1

1 +Tan2x =sec2x

(sec2x -1)(sec2x)

= sec4x -sec2x